Keith has $500 in a savings account at the beginning of the summer. He wants to have at least $200 in the account by the end of the summer. He withdraws $25 each week for food, clothes, and movie tickets.
Write an inequality that represents Keith’s situation.
How many weeks can Keith withdraw money from his account?

Answer by Theo(13342)

the inequality is 500 - 25x >= 200

this insures that he will have at least 200 at the end of the summer.

subtract 200 from both sides of that inequality and add 25x to both sides of that inequality to get 500 - 200 >= 25x

simplify to get 300 >= 25x

divide both sides of that equation by 25 to get 300 / 25 >= x

simplify to get 12 >= x

12 >= x means x <= 12.

when x is smaller than or equal to 12, he will be guaranteed to have at least 200 in the account at the end of the summer.

when x = 12, what is left in the account is 500 - 25 * 12 = 200.

when x = 11, what is left in the account is 500 - 25 * 11 = 225.

when x = 13, what is left in the account is 500 - 25 * 13 = 175.

the maximum number of weeks he can withdraw money from his account is 12.